Find the x and y intercepts of each linear equation. Be sure to write your answers as ordered pairs.
A. y = 3x - 6
B. 6y = -x + 2

Answer by checkley77(12569)

A. y=3x-6 (0,-6)
0=3X+6
3X=-6
X=-6/3
X=-2 OR (-2,0)
B. 6y=-x+2
Y=-X/6+2/6 (0,2/6)
0=-X/6+2/6
-X/6=-2/6
-X=-2
X=2 OR (2,0)